OpenStudy (anonymous):

need help finding 2 solutions for this equation.

OpenStudy (anonymous):

\[y ^{2\over3}+5y ^{1\over3}-104=0\]

OpenStudy (dumbcow):

treat it just like a quadratic equation since (y^1/3)^2 = y^2/3 factor , what factors of -104 add up to 5

OpenStudy (anonymous):

i couldn't think of anything that does.

OpenStudy (dumbcow):

start with the obvious 52*2 then take half of 52 and double 2 26*4 13*8

OpenStudy (anonymous):

-13 and 8 work I guess

OpenStudy (dumbcow):

check the signs, they need to add up to positive 5

OpenStudy (anonymous):

13 and -8

OpenStudy (dumbcow):

ok so now (y^1/3 +13)(y^1/3 - 8) = 0 solve each part for y

OpenStudy (anonymous):

so 13- 1/3 and 8-1/3? Sorry, I know this is way more simple than I'm making it.

OpenStudy (anonymous):

i just don't know if it will end up not being a whole number this way? and what that would be exactly

OpenStudy (dumbcow):

thats ok y^1/3 = -13 to get y by itself i have to take each side to the 3rd power this way (y^1/3)^3 = y y = -13^3 also y = 8^3

OpenStudy (anonymous):

oh ok. so the answers can be to the third power like that?

OpenStudy (anonymous):

my online assignment wont let me type the answer to the third power. Do I solve more from there?

OpenStudy (dumbcow):

put it in your calculator to get a whole number

OpenStudy (anonymous):

one of them is 28561 and the other is4096..? kinda big numbers.

OpenStudy (dumbcow):

8^3 = 8*8*8 = 512 -13^3 = -13*-13*-13 = -2197

OpenStudy (anonymous):

ohho k, i didn't know the answers could be that large.

OpenStudy (dumbcow):

only because of the fractions in the exponents